# Query-Efficient Zeroth-Order Algorithms for Nonconvex Optimization

### v1

@@ 2.1  Assumptions and Stationarity Measure
The assumption on the lower boundedness of $\Phi(x)$ is equivalent to assume that $h(x)$ is lower bounded for any $x\in\mathbb{R}^{d_{x}}$ satisfying $c_{j}(x)\leq 0,\forall j\in\mathcal{J}$.
@@
In our problem (2), while $\mathcal{Y}$ serves as the feasible set of Lagrange multipliers that are inherently unbounded, the boundedness of the optimal dual set has been justified in Nedić and Ozdaglar, (2009) under the Slater condition. Therefore, this assumption is commonly imposed in existing work (Liu et al., 2020b, ; Xu et al.,, 2023), and we can construct a bounded set containing the optimal dual variables to replace $\{y\in\mathbb{R}^{d_{y}}|y\geq 0\}$ in our method.
@@
$f(x,y)$ is differentiable and Lipschitz continuous, i.e., for any $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$, we have $\|\nabla_{x}f(x,y)\|\leq\Lambda$ and $\|\nabla_{y}f(x,y)\|\leq\Lambda$ for some $\Lambda>0$.
@@
$f(x,y)$ is $L$-smooth in $x$ and $y$, i.e., there exist some $L\geq 0$ satisfying $\|\nabla f(x_{1},y_{1})-\nabla f(x_{2},y_{2})\|\leq L\left(\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|\right)$ for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$, and $y_{1},y_{2}\in\mathcal{Y}$.
Assumptions 2.2-2.3 impose the Lipschitz continuity on $f(x,y)$ and its gradients, which are standard in the literature of both first-order and zeroth-order optimization (Nedić and Ozdaglar,, 2009; Ghadimi and Lan,, 2013; Zhou et al.,, 2025). Similarly, we can also equivalently impose Lipschitz continuity on $h(x),\;c_{j}(x),\forall j\in\mathcal{J}$ and their gradients to replace Assumptions 2.2 and 2.3.
For min-max problems, a widely adopted stationarity measure is the proximal gradient for first-order and zeroth-order nonconvex optimization (Lin et al.,, 2020; Liu et al., 2020b, ; Xu et al.,, 2023):
@@
where $\beta$ is the step size for dual updates. A point $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is a first-order stationary point of (2) if $\|\mathfrak{g}(x,y)\|=0$. We also introduce another notion of stationarity measure. The problem (2) is equivalent to minimizing the function $\Phi(x)=\max_{y\in\mathcal{Y}}f(x,y)$ over $\mathbb{R}^{d_{x}}$. The norm of $\nabla\Phi(x)$ is an appropriate stationarity measure for nonconvex optimization when $\Phi(x)$ is differentiable (Wang et al.,, 2023). However, $\Phi(x)$ may fail to be differentiable even if $f(x,y)$ is concave in $y$, as the maximum may not be uniquely attained. Alternatively, we define the Moreau envelope of $\Phi(x)$ for any $\lambda>0$ as
@@
The Moreau envelope $\Phi_{1/2L}(x)$ with parameter $\frac{1}{2L}$ and $\nabla\Phi_{1/2L}(x)$ are both well-defined because $\Phi(u)+L\|u-x\|^{2}$ is strongly convex in $u$ given $x$. Furthermore, $\Phi_{1/2L}(x)$ is differentiable and smooth in $x$. A point $x\in\mathbb{R}^{d_{x}}$ is a stationary point of $\Phi$ if $\|\nabla\Phi_{1/2L}(x)\|=0$. This stationarity measure is also widely used in nonconvex-concave settings (Mahdavinia et al.,, 2022; Davis and Drusvyatskiy,, 2019). As shown in Lin et al., (2020), the computational overhead of transferring this notion to the one measured by $\|\mathfrak{g}(x,y)\|$ is negligible compared to the overall query complexity. Therefore, we define our stationarity measure as follows.
@@
Let $M(x,y)=\min\left\{\|\mathfrak{g}(x,y)\|,\|\nabla\Phi_{1/2L}(x)\|\right\}$ for some $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$. We say a point $(x,y)$ is an $\epsilon$-stationary point of (2) if $M(x,y)\leq\epsilon$.
@@ 2.2  Zeroth-Order Gradient Estimation
where $g(x;r,z)\in\mathbb{R}^{d_{x}}$ is the estimated gradient for $\nabla h(x)$. Here, $r>0$ is the smoothing radius and $z\in\mathbb{R}^{d_{x}}$ is a random perturbation vector typically sampled from the Gaussian distribution $\mathcal{N}(0,I_{d_{x}})$ or the uniform distribution on a sphere with radius $\sqrt{d_{x}}$ (Nesterov and Spokoiny,, 2017; Duchi et al.,, 2015). When $h$ is smooth, this estimator enjoys a bias bounded by the smoothing radius (Malik et al.,, 2020). In contrast, CGE adds perturbation to each dimension separately and applies the finite difference of function values to construct a full gradient. The CGE is defined as
@@
where $g(x;r,\{e_{i}\}_{i=1}^{d_{x}})\in\mathbb{R}^{d_{x}}$ approximates $\nabla h(x)$, and $e_{i}\in\mathbb{R}^{d_{x}}$ is the unit vector with only the $i$th entry being 1. Let $g_{i}(x;r,e_{i})$ denote the $i$th entry of $g(x;r,\{e_{i}\}_{i=1}^{d_{x}})$. Similarly, the bias of CGE is also negligible given a small smoothing radius (Berahas et al.,, 2022).
Dilemma of Trading off Single-Step and Overall Query Complexities.
Both RGE and CGE have biases bounded by the smoothing radius $r$. RGE in (3) is efficient for a single step and only requires two function values to construct a gradient. Its variance approximately takes the form $\mathcal{O}(d)\|\nabla h(x)\|^{2}$ (Liu et al., 2020a, ). In unconstrained problems, we have $\|\nabla h(x^{*})\|=0$ for any optimal solution $x^{*}$. Therefore, the variance is negligible as the iterates approach the optimal solution, which allows RGE-based algorithms to mimic their first-order counterparts and achieve similar convergence results. However, in constrained problems, the above property does not hold, as the gradient $\nabla h(x^{*})$ may not be zero. The large variance of RGE leads to worse overall query complexities in constrained problems (Nguyen and Balasubramanian,, 2023). In contrast, the variance of CGE is controlled by the order of $\mathcal{O}(r^{2})$ and is negligible with small $r$. Therefore, CGE-based algorithms generally enjoy the state-of-the-art overall query complexity bounds (Xu et al.,, 2024; Zhou et al.,, 2025). However, CGE requires $\mathcal{O}(d)$ function values to construct a full gradient, which is inefficient for large $d$. As a result, achieving efficiency in both aspects has yet to be addressed.
@@ 3.1  Algorithm Design
We propose the zeroth-order block gradient descent ascent (ZOB-GDA) algorithm as presented in Algorithm 1 for solving (2). Our algorithm follows the main steps of standard GDA, which perform the gradient descent in $x$ and gradient ascent in $y$. However, ZOB-GDA differs from the conventional zeroth-order GDA method introduced in Liu et al., 2020b , which employs the RGE in (3). Instead of estimating a full gradient, ZOB-GDA randomly selects a block of dimensions to estimate the partial gradients and performs block descent ascent at each step. Specifically, for the $k$th iterate $(x_{k},y_{k})$, we randomly sample a block of dimensions, denoted by a set $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ ($b$ is an integer ranging from 1 to $d_{x}$). Then, we apply BCGE to estimate the block gradient of $f(x_{k},y_{k})$:
@@
where we denote $G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})\in\mathbb{R}^{d_{x}}$ as the vector with the entries of $\mathcal{I}_{k}$ being estimated and other entries being 0. We can apply different smoothing radii $r_{k}$ for each iteration. For simplicity, we denote $G_{x}(x_{k},y_{k})=G_{x}^{\mathcal{I}}(x_{k},y_{k})$ when $\mathcal{I}=[d_{x}]$. For the update of the dual variable $y$, we have the partial gradient $\nabla_{y}f(x_{k},y_{k})=c(x_{k})$. No additional queries are required as $c(x_{k})$ has been observed when computing $G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})$.
Specifically, when the block size satisfies $b=1$, the primal update resembles a coordinate update. When $b=d_{x}$, the primal update uses a full gradient and resembles the primal update in traditional CGE-based algorithms (Xu et al.,, 2024; Zhou et al.,, 2025). Moreover, we can control the single-step efficiency by adjusting the block size $b$. It will be shown that the choice of $b$ does not affect the overall query complexity of ZOB-GDA.
Unlike existing ZO literature that applies coordinate/block updates, we first generalize them within the GDA framework to address non-analytical constraints, whose dynamics and convergence analysis are significantly more complicated due to the coupling of primal and dual steps.
@@
3:  Randomly sample $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ and update $x_{k}$ by
@@ 3.2  Convergence Results of ZOB-GDA
Lemma 3.1 and its extended versions have appeared in existing literature (Lian et al.,, 2016; Berahas et al.,, 2022; Jin et al.,, 2023), thus, we omit its proof here. Lemma 3.1 demonstrates that $g_{i}(x;r,e_{i})$ is a good partial gradient estimator, in the sense that both its bias and variance can be effectively controlled by the smoothing radius and $L$. This error bound plays a fundamental role and will be frequently used in our theoretical analysis.
Let $N=\frac{d_{x}}{b}$, $R_{y}=\|\overline{y}\|$ and $\Lambda_{0}=\Lambda+\frac{\sqrt{b}L}{2}\cdot\sup_{k}\{r_{k}\}$. Then, we present the following Theorem to characterize the convergence of ZOB-GDA. Its proof is provided in Appendix A.
@@
Suppose Assumptions 2.1-2.3 hold. The sequence $\{(x_{k},y_{k})\}_{k=0}^{K}$ is generated by ZOB-GDA. The step sizes satisfy $0<\alpha,\beta\leq 1/L$, and the sequence of smoothing radii satisfies $\sum_{k=0}^{K}r_{k}^{2}<\frac{1}{b}$. Then, we have
$\min_{k\leq K-1}\mathbb{E}\left[\left\|\Phi_{1/2L}(x_{k})\right\|\right]\leq\mathcal{O}\left(\sqrt{\frac{N}{\alpha K}}\right)+\epsilon_{c},$
where $\epsilon_{c}=\big(16L\Lambda_{0}R_{y}\sqrt{2\alpha/\beta}+48\alpha L\Lambda_{0}^{2}\big)^{1/2}$.
The results in Theorem 3.1 imply that ZOB-GDA can converge within a fixed error at a convergence rate of $\mathcal{O}(\sqrt{N/\alpha K})$. The query complexity required to achieve $\min_{k\leq K-1}\mathbb{E}\left[M(x_{k},y_{k})\right]\leq\epsilon+\epsilon_{c}$ is $\mathcal{O}(d/\epsilon^{2})$. The fixed error results from the use of constant step sizes. Given the expression of $\epsilon_{c}$, one could adopt diminishing or small step sizes for $\alpha$ to eliminate the term $\epsilon_{c}$. The following corollary characterizes the exact convergence guarantee of ZOB-GDA.
@@
The derivation of Corollary 3.2 is straightforward by substituting the step size into the result in Theorem 3.1. The result in Corollary 3.2 shows that using two-time-scale step sizes for the updates of $x$ and $y$ can effectively eliminate the fixed error term $\epsilon_{c}$. This phenomenon aligns with the first-order GDA in nonconvex-concave settings (Lin et al.,, 2020). Building upon Corollary 3.2, we obtain the following corollary to establish the overall query complexity of ZOB-GDA.
@@
Suppose the conditions in Theorem 3.1 hold. Set $\alpha=\mathcal{O}\left(\epsilon^{4}\right)$ for any sufficiently small $\epsilon$. Then, the query complexity to achieve $\min_{k\leq K-1}\mathbb{E}\left[\left\|\nabla\Phi_{1/2L}(x_{k})\right\|\right]\leq\epsilon$ is $\mathcal{O}\left(\frac{d_{x}}{\epsilon^{6}}\right)$.
To the best of our knowledge, Corollary 3.3 establishes the first query complexity result for zeroth-order algorithms in the standard GDA framework for nonconvex-concave settings. Notably, this complexity differs from the first-order GDA by an additional factor $d_{x}$ (Lin et al.,, 2020), which is inherent to zeroth-order gradient estimation. Compared to the ZOAGP algorithm (Xu et al.,, 2024) with the query complexity bound $\mathcal{O}(d/\epsilon^{4})$, ZOB-GDA’s complexity bound seems worse due to the limitation of standard GDA framework. However, the single-step gradient estimation can be significantly more efficient for ZOB-GDA by using a small $b$. In the next section, we will leverage block updates with a variant of the GDA framework to design a new algorithm that achieves both the best-known overall query complexity and adjustable single-step efficiency.

### v2

@@ 2.1  Assumptions and Stationarity Measure
A sufficient condition for the lower boundedness of $\Phi(x)$ is that $h(x)$ is lower bounded for any $x\in\mathbb{R}^{d_{x}}$. In problem (2), the natural domain of Lagrange multipliers is $\mathbb{R}^{d_{y}}_{+}$. Nevertheless, restricting the dual iterates to a compact set $\mathcal{Y}$ that contains at least one optimal multiplier is standard in primal–dual analysis. Under Slater condition, the boundedness of the optimal dual set has been justified in Nedić and Ozdaglar (2009). This assumption is commonly imposed in the analysis of GDA-type algorithms and is essential for establishing finite-time convergence guarantees (Liu et al., 2020b; Xu et al., 2023).
@@
$f(x,y)$ is differentiable and Lipschitz continuous, i.e., $\forall(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$, we have $|f(x_{1},y_{1})-f(x_{2},y_{2})|\leq\Lambda\|(x_{1},y_{1})-(x_{2},y_{2})\|$ for some $\Lambda>0$.
@@
$f(x,y)$ is $L$-smooth in $x$ and $y$, i.e., there exist some $L>0$ satisfying $\|\nabla f(x_{1},y_{1})-\nabla f(x_{2},y_{2})\|\leq L\|(x_{1},y_{1})-(x_{2},y_{2})\|$ for any $x_{1},x_{2}\in\mathbb{R}^{d_{x}}$, and $y_{1},y_{2}\in\mathcal{Y}$.
Assumptions 2.2-2.3 impose the Lipschitz continuity on $f(x,y)$ and its gradients, which are standard in the literature of both first-order and zeroth-order optimization (Nedić and Ozdaglar, 2009; Ghadimi and Lan, 2013; Zhou et al., 2025). Similarly, we can also equivalently impose Lipschitz continuity on $h(x),\;c_{j}(x),\forall j\in\mathcal{J}$ and their gradients to replace Assumptions 2.2 and 2.3.
For min-max problems, a widely adopted stationarity measure is the proximal gradient for first/zeroth-order nonconvex optimization (Lin et al., 2020; Liu et al., 2020b; Xu et al., 2023):
@@
where $\beta$ is the step size for dual updates. A point $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is a first-order stationary point of (2) if $\|\mathfrak{g}(x,y)\|=0$.
We also introduce another notion of stationarity measure. The problem (2) is equivalent to minimizing the function $\Phi(x)=\max_{y\in\mathcal{Y}}f(x,y)$ over $\mathbb{R}^{d_{x}}$. The norm of $\nabla\Phi(x)$ is an appropriate stationarity measure for nonconvex optimization when $\Phi(x)$ is differentiable (Wang et al., 2023). However, $\Phi(x)$ may fail to be differentiable even if $f(x,y)$ is concave in $y$, as the maximum may not be uniquely attained. Alternatively, we define the Moreau envelope of $\Phi(x)$ for any $\lambda>0$ as
@@
Under Assumption 2.3, the Moreau envelope $\Phi_{1/2L}(x)$ with parameter $\lambda=\frac{1}{2L}$ and $\nabla\Phi_{1/2L}(x)$ are both well-defined because $\Phi(u)+L\|u-x\|^{2}$ is strongly convex in $u$ given $x$. Furthermore, $\Phi_{1/2L}(x)$ is differentiable and smooth in $x$. A point $x\in\mathbb{R}^{d_{x}}$ is a stationary point of $\Phi$ if $\|\nabla\Phi_{1/2L}(x)\|=0$. This stationarity measure is also widely used in nonconvex-concave settings (Mahdavinia et al., 2022; Davis and Drusvyatskiy, 2019). As shown in Lin et al. (2020), the computational overhead of transferring this notion to the one measured by $\|\mathfrak{g}(x,y)\|$ is negligible relative to the overall query complexity. With this minor distinction, we define our stationarity notion as follows.
@@
We say a point $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is an $\epsilon$-stationary point of (2) if $\|\mathfrak{g}(x,y)\|\leq\epsilon$ or $\|\nabla\Phi_{1/2L}(x)\|\leq\epsilon$.
In this study, we therefore adopt different metrics of stationarity depending on the analysis techniques, since both metrics are commonly applied.
@@ 2.2  Zeroth-Order Gradient Estimation
where $g(x;r,z)\in\mathbb{R}^{d_{x}}$ is the estimated gradient for $\nabla h(x)$. Here, $r>0$ is the smoothing radius and $z\in\mathbb{R}^{d_{x}}$ is a random perturbation vector typically sampled from the Gaussian distribution $\mathcal{N}(0,I_{d_{x}})$ or the uniform distribution on a sphere with radius $\sqrt{d_{x}}$ (Nesterov and Spokoiny, 2017; Duchi et al., 2015). In contrast, CGE adds perturbation to each dimension separately and applies the finite difference of function values to construct a full gradient. The CGE is defined as
@@
where $g(x;r,\{e_{i}\}_{i=1}^{d_{x}})\in\mathbb{R}^{d_{x}}$ approximates $\nabla h(x)$, and $e_{i}\in\mathbb{R}^{d_{x}}$ is the unit vector with only the $i$th entry being 1. Let $g_{i}(x;r,e_{i})$ denote the $i$th entry of $g(x;r,\{e_{i}\}_{i=1}^{d_{x}})$.
Challenge of Trading off Single-Step and Overall Query Complexities.
It has been shown that the biases of both (3) and (4) are bounded by $\mathcal{O}(\sqrt{d_{x}}r)$ (Berahas et al., 2022), which means the biases are negligible when a small $r$ is applied. RGE in (3) is efficient for a single step and only requires two function values to construct a gradient. Its variance approximately takes the form $\mathcal{O}(d_{x}\|\nabla h(x)\|^{2}+d_{x}^{3}r^{2})$ (Liu et al., 2020a). In unconstrained problems, we have $\|\nabla h(x^{*})\|=0$ for any optimal solution $x^{*}$. Therefore, the variance of RGE shrinks as the iterations approach the optimal solution, which allows RGE-based algorithms to mimic their first-order counterparts and achieve similar convergence results. However, in constrained problems, the above property does not hold, as the gradient $\nabla h(x^{*})$ may not be zero. The large variance of RGE leads to worse overall query complexity bounds in constrained problems (Jin et al., 2026). In contrast, the variance of CGE is controlled by the order of $\mathcal{O}(d_{x}r^{2})$ and is negligible with small $r$. Therefore, CGE-based algorithms generally enjoy the state-of-the-art overall query complexity bounds (Xu et al., 2024). However, CGE requires $\mathcal{O}(d_{x})$ function values to construct a full gradient, which is inefficient for a large $d_{x}$.
As a result, achieving efficiency in both aspects has not been well addressed. Below, we will show that BCGE is a promising answer to this open question when applied to solve (1).
@@ 3.1  Algorithm Design
We propose the zeroth-order block gradient descent ascent (ZOB-GDA) algorithm as presented in Algorithm 1 for solving (2). Our algorithm follows the main steps of standard GDA, which perform the gradient descent in $x$ and gradient ascent in $y$. However, ZOB-GDA differs from the conventional zeroth-order GDA method introduced in Liu et al. (2020b), which employs the RGE in (3). Instead of estimating a full gradient, ZOB-GDA randomly selects a block of dimensions to estimate the partial gradients and performs block descent ascent at each step. Specifically, for the $k$th iterate $(x_{k},y_{k})$, we uniformly sample a block of dimensions, denoted by a set $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ ($b$ is the block size ranging from 1 to $d_{x}$). Then, we estimate the BCGE of $f(x_{k},y_{k})$ by:
@@
where we denote $G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})\in\mathbb{R}^{d_{x}}$ as the vector with the entries of $\mathcal{I}_{k}$ being estimated and other entries being 0. We can apply different smoothing radii $r_{k}$ for each iteration. For simplicity, we denote $G_{x}(x_{k},y_{k})=G_{x}^{\mathcal{I}}(x_{k},y_{k})$ when $\mathcal{I}=[d_{x}]$. For the update of the dual variable $y$, we have the partial gradient $\nabla_{y}f(x_{k},y_{k})=c(x_{k})$. Therefore, no additional queries are required as $c(x_{k})$ has been observed when computing $G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})$. Therefore, $b+1$ queries are required for each iteration.
Specifically, when the block size satisfies $b=1$, the primal update resembles a coordinate update. When $b=d_{x}$, the primal update uses a full gradient and resembles the primal update in traditional CGE-based algorithms (Xu et al., 2024; Zhou et al., 2025). Moreover, we can control the single-step efficiency by adjusting the block size $b$. It will be shown that the choice of $b$ does not affect the theoretical query complexity bound of ZOB-GDA.
Unlike existing ZO literature that applies coordinate/block updates, we first generalize them in the GDA framework to address non-analytical constraints, whose dynamics and convergence analysis are significantly more complicated due to the coupling of primal and dual steps.
@@
3:  Uniformly sample $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ and update $x_{k}$ by
@@ 3.2  Convergence Results of ZOB-GDA
Lemma 3.1 and its extended versions have appeared in existing literature (Lian et al., 2016; Berahas et al., 2022; Jin et al., 2026), thus, we omit its proof here. Lemma 3.1 also implies that $\left|\nabla_{i}h(x)-g_{i}(x;r,e_{i})\right|^{2}\leq\frac{1}{4}L^{2}r^{2}$. It demonstrates that $g_{i}(x;r,e_{i})$ is a good partial gradient estimator, in the sense that both its bias and variance can be effectively controlled by the smoothing radius. This error bound plays a fundamental role and will be frequently used in our theoretical analysis.
Let $N=\frac{d_{x}}{b}$, $D_{y}=\|\overline{y}\|$ and $\Lambda_{0}=\Lambda+\frac{\sqrt{b}L}{2}\cdot\sup_{k}\{r_{k}\}$. Then, we present the following Theorem to characterize the convergence of ZOB-GDA in the sense of $\left\|\nabla\Phi_{1/2L}(x_{k})\right\|$.
@@
Suppose Assumptions 2.1-2.3 hold. The sequence $\{(x_{k},y_{k})\}_{k=0}^{K}$ is generated by ZOB-GDA. The step sizes satisfy $0<\alpha,\beta\leq 1/L$, and the sequence of smoothing radii satisfies $\sum_{k=0}^{K}r_{k}^{2}\leq\frac{1}{b}$. Then, we have
$\min_{k\leq K-1}\mathbb{E}\left[\left\|\nabla\Phi_{1/2L}(x_{k})\right\|\right]\leq\mathcal{O}\left(\sqrt{\frac{N}{\alpha K}}\right)+\epsilon_{c},$
where $\epsilon_{c}=\big(16L\Lambda_{0}D_{y}\sqrt{2\alpha/\beta}+32\alpha L\Lambda_{0}^{2}\big)^{1/2}$.
The results in Theorem 3.1 imply that ZOB-GDA can converge to an $\epsilon_{c}$-stationary point at a convergence rate of $\mathcal{O}(\sqrt{N/\alpha K})$. The fixed error $\epsilon_{c}$ results from the use of constant step sizes for both primal and dual steps, which is also observed for the first-order GDA (Nedić and Ozdaglar, 2009). We provide the proof of Theorem 3.1 in Appendix A.
Given the expression of $\epsilon_{c}$, one could adopt diminishing or small step sizes for $\alpha$ to eliminate the fixed term $\epsilon_{c}$. The following corollary characterizes the convergence guarantee of ZOB-GDA.
@@
The derivation of Corollary 3.2 is straightforward by substituting the step size into the result in Theorem 3.1. Since the step size $\alpha$ is dependent on the maximum steps $K$, we can apply diminishing settings for $\alpha$ in practice. Corollary 3.2 shows that using two-time-scale step sizes for the updates of $x$ and $y$ can effectively eliminate the fixed error term $\epsilon_{c}$. Building upon Corollary 3.2, we obtain the following corollary to establish the overall query complexity bound of ZOB-GDA.
@@
Suppose the conditions in Theorem 3.1 hold. Set $\alpha=\mathcal{O}\left(\epsilon^{4}\right)$ for any sufficiently small $\epsilon>0$. Then, the query complexity to achieve $\min_{k\leq K-1}\mathbb{E}\left[\left\|\nabla\Phi_{1/2L}(x_{k})\right\|\right]\leq\epsilon$ is $\mathcal{O}\left(\frac{d_{x}}{\epsilon^{6}}\right)$.
To the best of our knowledge, Corollary 3.3 establishes the first query complexity bound for zeroth-order algorithms in the standard GDA framework for nonconvex-concave settings. Notably, this complexity bound differs from the first-order GDA by an additional factor $d_{x}$ (Lin et al., 2020), which is inherent to zeroth-order gradient estimation and occurs commonly in zeroth-order optimization (Liu et al., 2020a). Compared to the ZOAGP algorithm (Xu et al., 2024) with the query complexity bound $\mathcal{O}(d/\epsilon^{4})$, ZOB-GDA’s complexity bound seems worse due to the limitation of standard GDA frameworks. However, the single-step gradient estimation can be significantly more efficient for ZOB-GDA by using a small block size. In the next section, we will leverage block updates with a variant of the GDA framework to design a new algorithm that achieves both the best-known overall query complexity bound and adjustable single-step efficiency.